OpenStudy (anonymous):

Solve for x: -4(x-3)=16

OpenStudy (anonymous):

x=-4/3

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

-4x +12 = 16 -4x = 4 x = (4/-4) x = -1